Status Report

NASA Space Station On-Orbit Status 9 Dec 2003

By SpaceRef Editor
December 9, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 9 Dec 2003

All ISS systems continue to function nominally, except as noted previously or below.

After a number of regular maintenance chores, the crew started out on the major TVIS treadmill IFM (in-flight maintenance), for internal inspection, modification and replacements. They completed almost all of the tasks planned for today. Finishing up and reassembly of the treadmill is scheduled for tomorrow.[Today’s 4-hr. job involved: (a) removal of the big machine from the Service Module (SM) floor pit, (b) resetting the “Enable” DIP switch of the horizontal roll-stabilizing gyrodyne to uncouple the yaw and pitch stabilizers for independent function, and (c) removal & opening the gyro housing for inspection and photography. Removal of clamp bracket assemblies at the two aft top corners and replacement of the four frayed wire rope loops at those locations plus retightening the screws (four per corner), are scheduled for tomorrow. On the gyro, two discoveries were made: (a) deposits of “fine black powder” in some places, and (b) an apparently broken top bearing of the horizontal gyro disk (as demonstrated by the crew, it resisted spinning, while the bottom bearing spun freely).]

Before the IFM, CDR Foale downloaded the accumulated TVIS, RED expander and CEVIS cycle exercise files to the MEC (medical equipment computer) for subsequent downlink via S-band comm on OCA.

FE Kaleri continued the latest maintenance round on the Russian segment (RS) ventilation system, today replacing the four PF1-4 dust collector filters in the SM with new cartridges (last time done: 11/5).

Kaleri terminated the bake-out cycle on filter bed 1 of the SM’s harmful impurities (atmospheric trace contaminants) removal unit (BMP), moding the channel back to Purify. Later, he initiated regeneration on filter channel 2.[Regeneration of the air purifier filter beds is repeated every 20 days. Each bakeout to space vacuum takes about 24 hours.]

The SO continued work on the MSG PFMI (Microgravity Science Glovebox/Pore Formation & Mobility Investigation) experiment to support the ongoing troubleshooting.[After a circuit breaker interrupted the experiment run yesterday in an early stage, it was aborted. Today, Mike activated the MSG to allow the ground to downlink the data, along with digital photos of the current configuration, for analysis by specialists. Afterwards, MSG was deactivated.]

Mike also completed the deployment of the EHS (environmental health systems) acoustic dosimeters, rescheduled today from yesterday.[Today’s task applied only to the crew-worn dosimeters; they were donned at 1:00am EST this morning for first read-off and recording tonight at 4:10pm. Second read-off and recording will follow tomorrow morning at 1:00am, to conclude the required 24-hr. period. The instruments will then be stowed. Surgeon will reschedule the deployment of the single static acoustic dosimeter during next week.]

Sasha completed the daily RS life support systems (SOZh) maintenance, including toilet facility, food containers, water containers and solid waste containers, and prepared the daily IMS update file.

Yesterday the Russian VELO cycle ergometer (VB-3) sustained a failure, apparently a recurrence of the control module problem seen previously (10/31). The cycle didn’t work for a time until the crew successfully recovered it. It is now functioning again. [The recent troubleshooting by Kaleri failed to identify the cause.]

After the recent (11/27) starboard & zenith video inspection of the SM exterior with MSS (mobile service system) cameras, the next Robotics proficiency operations is scheduled for 12/11 (Thursday). The associated DOUG (dynamic operations ubiquitous graphics) application will be set up and reviewed tomorrow.[The crew will use Canadarm2/SSRMS (Space Station Remote Manipulator System) for inspecting the SM portside to check for any indications of the cause of the noise heard on 11/26. The survey will also cover the P6 4-bar linkage shoulder bolts to determine if the bolts have backed out, as well as provide general views of the zenith (top) side of the ISS, including the FGB and the Z1 and P6 truss segments, to look for anything anomalous.]

Starting at 1:30am EST tomorrow morning and running for seven hours, MCC-Houston and its support room in Moscow (HSR) will conduct another periodic BCC (backup control center) dry-run in test mode, with no involvement of the ISS crew or vehicle. This checks up on preparedness for a contingency ISS flight control handover from Houston to Moscow.[BCC utilizes the HSR command server and telemetry server. In test mode, HSR (as BCC) will send a single not-for-uplink PPCP (preplanned command packet) to MCC-M/TsUP, then cancels it after successful transmission is confirmed. MCC-M plays back previously recorded ACT (American contingency telemetry), with HSR confirming successful receipt. TsUP commanding to the Russian segment (RS) via U.S. S-band as well as MCC-H-to-TsUP flow of status telemetry are also not affected, but MCC-H has no real-time telemetry and no command capability via Russian ground stations (RGS) during the dry-run.]

Prioritizing of upmass manifested on Progress 13P (launch NET 1/28/04) continues. Two new TVIS gyro bearings, along with other already identified TVIS parts, will be sent to the station. Other U.S. hardware approved for 13P include two IMAKs (ISS Medical Accessory Packs) with fresh medications, two advanced SchRED (Schwinn resistive exercise device) canisters, and a new MCA (major constituents analyzer).

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, wereKarachi, Pakistan (looking to nadir and a touch left. No views of this important South Asian port exist at the necessary resolution. The coastline of the Indus River delta, where Karachi is situated, has advanced several miles seaward since Alexander the Great passed through the region in 325 BC), Smoke, smog, Pakistan & India (Dynamic event: The smog event in northern India continues [visibility in Delhi is less than 3/4 mile]. Looking right for aerosols in the Ganges basin), Swampland mapping, SW Sudan (views looking from nadir and off left of track up to 2 degrees were requested to document poorly known landscapes, especially highly mobile stream courses, W of the Nile. A complex of inland deltas appears to be present. These landscapes are generating scientific interest since the scale of inland deltas has been unappreciated by geologists), Atlanta, Georgia (looking a touch right. Atlanta is spread through several well-wooded valleys and its full extent can be difficult to determine), Washington, D.C. (looking a touch right. Also trying for obliques of the widespread snowfall that tied up the NE recently), Philadelphia, PA (looking a touch right), andJarvis Island, Pacific (this trapezoid island is about 2 miles long and completely surrounded by a coral reef. Detail of the coral reef is requested for the reef atlas).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:00pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 18A. Vozdukh CO2 scrubber is On (in Manual Mode 5/3). U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. MCA (major constituents analyzer) is off (in Life Extending Mode). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Regeneration mode. RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 26.7; ppO2 (mmHg) — 155.6; ppCO2 (mmHg) — 4.9.
  • SM Transfer Compartment: Pressure (mmHg) — 761; temperature (deg C) — 20.0.
  • FGB Cabin: Pressure (mmHg) — 756; temperature (deg C) — 22.7.
  • Node: Pressure (mmHg) — 753.43; temperature (deg C) — 23.7 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 755.62; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 755.72; temperature (deg C) — 27.4; shell heater temp (deg C) — 24.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1: Shell heater temp (deg C) — 22.4
  • PMA-2: Shell heater temp (deg C) — 13.5

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Directed Position (non-suntracking).
  • SM batteries: Battery #8, formerly known as #7, is still disconnected in slot #8 for troubleshooting, off-line; all other batteries (7) are in “Partial Charge” mode.
  • FGB batteries: Battery #5 is off (capacity restoration mode, ROM); battery #3 is in “Cycle Mode” ; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available:3703 kg (8164 lb) as of 12/4/03[SM (755) + FGB (2596) + Progress M (352) + Progress M-1 (0)]. (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed” : z-axis in local vertical, y-axis in velocity vector [yaw: 90.2 deg, pitch: -10.8 deg, roll: -3 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime, IAC-1 is suspect).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings.
  • MT: latched and mated at WS4.
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit (as of this morning, 7:27am EST [= epoch]):

  • Mean altitude — 371.8 km
  • Apogee — 376.0 km
  • Perigee — 367.7 km
  • Period — 92.0 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0006109
  • Orbits per 24-hr. day — 15.65
  • Mean altitude loss in last 24 hours — 100 m
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 28841
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.